HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH RESOLUTION SPECTROSCOPY OF RELATIVELY LARGE RADICAL SPECIES IN THE SUB-MILLIMETER WAVE RANGE USING ZEEMAN AND FARADAY ROTATION ACQUISITIONS. FIRSTS RESULTS

ON VINOXY, ACETYL, AND ACETONYL RADICALS.

\underline{R. CHAHBAZIAN} , \underline{L. JUPPET} \underline{M.-A. MARTIN-DRUMEL} , \underline{O. PIRALI} , Université Paris-Saclay, CNRS, Institut des Sciences Moléculaires d’Orsay, 91400

Orsay, France

Molecular radicals are key species for chemical reactions occurring in combustion processes as well as in atmospherical and astrophysical environments^{1}. However, the characterization of “large” radicals (defined here as containing more than 5 atoms) using high-resolution gas phase absorption spectroscopic techniques remains tedious and requires the development of specific experimental setups to increase the sensitivity of the spectrometers. To this end, we associated a frequency multiplication chain spectrometer, allowing high-resolution spectral acquisition in the 75-900 GHz range, with a relatively specific technique of radical synthesis based on H abstraction on a given precursor by reaction with F atoms in a flow cell^{2}^{,}^{3}. Since paramagnetic species (e.g., radicals) are sensitive to the magnetic field, an alternative magnetic field was generated in the cell allowing the use of Zeeman Modulation^{4} (ZM) as well as Faraday Rotation Modulation^{5} (FRM) techniques to record broadband and “radical-only” spectra. Using these techniques, the spectra of the vinoxy radical (CH_{2}CHO), the acetyl radical (CH_{3}CO), and the acetonyl radical (CH_{3}COCH_{2}) have been recorded. I will present these instrumental developments as well as the associated preliminary spectroscopic results.

1Hirota, Eizi, Springer Series in Chemical Physics, 40 (1985)

2G. Inoue and H. Akimoto, J. Chem. Phys. 74 425 (1981)

3O. Chitarra, O. Pirali, J.-T. Spaniol, T. S. Hearne, J.-C. Loison, J. F. Stanton, and M.-A. Martin-Drumel, The Journal of Physical Chemistry A, 126 7502 (2022)

4M.Elhanine, R.Farrenq and G.Guelachvili, Mikrochimica Acta, 95 265 (1988)

5G. Litfin, C.-R. Pollock, R.-F. Curl and F.-K. Tittel, The Journal of Chemical Physics, 72 6602 (1980)

p-number: p054  Submitted on Thu, 08 Jun 2023 16:46:55 +0200